OpenStudy (anonymous):

The region bounded by the given curves is rotated about the specified axis. Find the volume V of the resulting solid by any method. x = (y − 2)2, x = 1; about y = 1

OpenStudy (anonymous):

This is the graph. The region enclosed has to be rotated around the black line

OpenStudy (anonymous):

Should I use shell or washer on this?

OpenStudy (anonymous):

what's the prob? you should be a pro at this now!

OpenStudy (anonymous):

Haha, I know, I always feel confident, set them up, work them out but then when I turn it in, I always get them wrong.

OpenStudy (anonymous):

I initially set it up using shell. I had r(y)=y and h(y)=(y-2)^2

OpenStudy (anonymous):

r = (y-1) since you're rotating around the line y = 1. limits are 1 to 3. h = (1-(y-2)^2)

OpenStudy (anonymous):

Is r = y-1 ?

OpenStudy (anonymous):

Ahhh, yes, okay. I had a realization right before you said that, lol

OpenStudy (anonymous):

good job!

OpenStudy (anonymous):

Yeah. :b But when I worked it out that time I got a negative answer ? 0.0

OpenStudy (anonymous):

hold on a sec... is this what you got for your integral?